# Deep Semantic Inference over the Air: An Efficient Task-Oriented Communication System

### v1

@@ Abstract
Empowered by deep learning, semantic communication marks a paradigm shift from transmitting raw data to conveying task-relevant meaning, enabling more efficient and intelligent wireless systems. In this study, we explore a deep learning-based task-oriented communication framework that jointly considers classification performance, computational latency, and communication cost. We adopt ResNets-based models and evaluate them on the CIFAR-10 and CIFAR-100 datasets to simulate real-world classification tasks in wireless environments. We partition the model at various points to simulate split inference across a wireless channel. By varying the split location and the size of the transmitted semantic feature vector, we systematically analyze the trade-offs between task accuracy and resource efficiency. Experimental results show that, with appropriate model partitioning and semantic feature compression, the system can retain over 85% of baseline accuracy while significantly reducing both computational load and communication overhead.
Index Terms:
Deep learning, task-oriented communication, wireless.
@@ I Introduction
Traditional wireless communication systems have been designed according to Shannon’s information theory, where the primary objective is to reliably transmit bits over noisy channels [1]. However, with the rapid rise of data-driven applications such as autonomous driving, real-time surveillance, and industrial automation, communication networks are now expected to support not only reliable data transfer, but also intelligent, task-specific decision-making. This paradigm shift has prompted a rethinking of communication system design, leading to increased interest in semantic and task-oriented communication [2]. Unlike conventional systems that aim to preserve all data uniformly, semantic communication focuses on conveying the essential information relevant to the underlying task, aligning more closely with how humans communicate efficiently and robustly in uncertain environments [3, 4, 5].
@@
Recent advances in machine learning, particularly deep learning, have significantly accelerated the development of the semantic communication systems. Deep neural networks (DNNs), such as convolutional neural networks (CNNs) for images or Transformers for language, can capture high-level abstractions. Semantic encoders and decoders can be co-trained with downstream inference tasks to extract and transmit only the most relevant features. For example, instead of sending the full-resolution image, the encoder may learn to transmit only task-critical semantic features (e.g., class labels, bounding boxes, or embeddings), greatly reducing communication cost without compromising task performance. This has further led to the task-oriented communication systems [6], where performance is measured not by bit error rate but by how well a specific task (e.g., classification, detection, or control) is completed. Such systems are particularly valuable in resource-constrained environments like wireless sensor networks or mobile edge computing, where devices operate with limited energy, bandwidth, and processing power.
@@
Despite promising progress, designing and deploying deep learning-based task-oriented communication systems remains a complex challenge. These systems must seamlessly integrate sensing, encoding, transmission, and inference into a unified pipeline. Decisions regarding model architectures, partitioning strategies (e.g., split computing between edge and cloud), and training objectives must be carefully aligned with system-level constraints such as latency, energy consumption, and bandwidth limitations [7, 8, 9]. While deep learning models can be trained offline using powerful infrastructure, task execution must often meet strict real-time requirements. Inference needs to be completed both accurately and timely, making it essential to balance task performance with efficient use of computational and communication resources at runtime. Furthermore, existing studies often lack a unified analysis across different deployment settings, motivating a more systematic investigation of model split strategies and resource trade-offs.
@@
In this paper, we aim to address these challenges by designing and evaluating a deep learning-based task-oriented communication framework that explicitly considers both task accuracy and latency. The latter includes both computation delay (from model inference) and communication delay (from transmitting semantic features). We focus on classification tasks, adopting 18-layer and 34-layer Residual Networks [10] as the backbone models and using the CIFAR-10 and CIFAR-100 datasets [11] as benchmarks. We explore various model partitioning strategies and evaluate how the dimension of transmitted semantic features affects the trade-offs between task accuracy, computational load, and communication cost.
@@
The remainder of this paper is organized as follows. Section II introduces the system architecture and formalizes task latency. Section III presents the design and implementation of the classification system. Section IV evaluates the system under various configurations and analyzes the trade-offs between accuracy and latency. Section V concludes the paper and discusses future directions.
@@ II Task Latency in Deep Learning-Based Task-Oriented Communication
  • Transmitter-side inference: The classifier is fully deployed at the transmitter (e.g., an edge device), and only the final classification result (e.g., a label) is sent. This minimizes communication overhead but places high computational demands on resource-constrained edge devices.
  • Receiver-side inference: The full input is transmitted to the receiver (e.g., a cloud server), which performs the classification. This reduces transmitter-side computation but incurs high communication cost.
@@
In this work, we adopt the third approach, i.e., split inference. We illustrate the system architecture in Fig. 1. Let $M_{t}$ and $M_{r}$ denote the parts of the DNN model deployed at the transmitter and receiver, respectively. Their inference times are denoted as $T_{M_{t}}$ and $T_{M_{r}}$. In general, the inference time of a model depends on both its complexity and the hardware it runs on. We quantify model complexity using floating-point operations (FLOPs), denoted by $F_{M}$ for the model $M$, which provides a hardware-agnostic measure of computational demand. The actual computation time of the model $M$ can be estimated as a function of its FLOP counts, where the function is determined by the hardware settings of the device that runs the model. Inside the compute-bound regime, the function is linear and monotone, so we estimate the computation time as:
@@
[…]
In subsequent sections, we design a task-oriented communication system for classification, and evaluate the trade-off between task latency and accuracy under different system configurations, focusing on the model partitioning strategy and the value of $N_{c}$. Our objective is to design a system that meets real-time task execution requirements while efficiently utilizing available computational and communication resources.
@@ III System Design for Task-Oriented Classification
Figure 2: Illustration of ResNet-34 architecture and a split-inference task-oriented network based on the ResNet-34 Vanilla model (unmodified baseline version). Specifically, the first three parts (input, conv1, and conv2_x) are considered in the transmitter (encoder) as pre-processing, while the rest of them are considered in the receiver (decoder) as post-processing. The compression and decompression modules ensure the post-processing input is consistent with the pre-processing output while preserving invariance and equivariance to translation. The semantic representation with $z$ dimension is transmitted from the encoder to the decoder over the AWGN channel. The FLOPs and proportion of each module are listed below.
@@ III-A Deep Learning Model for Classification
We adopt the 18-layer and the 34-layer Residual Networks (ResNets) as the backbone models for our task-oriented communication system. These widely used CNN architectures are known for their robustness, modularity, and strong performance across a broad range of image classification tasks.
@@
Due to their widespread adoption and adaptability across datasets and domains, ResNet-based architectures like ResNet-34 make our system design generalizable to a variety of classification tasks and application scenarios.
@@ III-C Wireless Channel Simulation with AWGN
where $\sigma^{2}$ denotes the noise variance and is controlled by the assigned signal-to-noise ratio (SNR) of the channel. The receiver obtains the noisy feature vector $\tilde{z}$ and restores it to its pre-compression size, then uses it as the input to $M_{r}$ to perform the remaining convolutional layers, if any, and finally the classification task. This setup enables us to assess the robustness and performance of the system under varying channel conditions.
To ensure a seamless interface between the encoder and decoder, the output dimension of $M_{t}$ should match the input dimension of $M_{r}$, i.e., both must operate on a vector of size $N_{c}$. Training and evaluation are implemented in an end-to-end manner within the widely adopted deep learning framework, PyTorch [12]. The source code is available at the GitHub repository: https://github.com/SanyeungWang/Deep-Semantic-Inference-over-the-Air.
@@ IV-A Datasets
CIFAR-10: This dataset [11] contains 60,000 color images of size $32\times 32$ across 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. Each class includes 6,000 images. The dataset is divided into five training batches and one test batch, each with 10,000 images. Each RGB image has a shape of $32\times 32\times 3$, and class labels are encoded as integers from 0 to 9. In our experiments, we use the built-in CIFAR-10 dataset provided by PyTorch.
CIFAR-100: Also introduced in [11], CIFAR-100 shares the same image format as CIFAR-10 but includes 100 classes with 600 images each. For every class, there are 500 training images and 100 test images, making the classification task more fine-grained. The dataset is split into a single training batch of 50,000 images and a test batch of 10,000 images. We use the built-in CIFAR-100 dataset from PyTorch.
@@ IV-C Comparisons With Different Settings
Based on the proposed task-oriented communication system, we first analyze performance under different $z$ dimensions and SNR levels, as shown in Fig. 3. We evaluate both ResNet-18 and ResNet-34, each split between conv_2 and conv_3 (referred to as SP-2), with one fractionally-strided convolution module (ConvTranspose2d) in the decompression module, and assuming unlimited computing power in the cloud device initially. Top-1 Accuracy refers to the comparison between the class with the highest probability and the ground-truth label. With CIFAR-10, as shown in Figs. 3(a) and 3(b), we observe that higher SNR leads to better performance, but accuracy approaches 0.9 even for low-dimensional $z$. This saturation effect makes CIFAR-10 less suitable for analyzing the impact of model architecture and compression settings. In contrast, CIFAR-100 (Figs. 3(c) and 3(d)) achieves lower overall accuracy due to its finer-grained classes and smaller inter-class variance. The performance gap between different SNR levels is also more pronounced, highlighting its sensitivity to AWGN and its suitability for studying the effects of semantic compression and noise. Therefore, we focus on CIFAR-100 in subsequent experiments.
Figure 3: Comparison of Top-1 Accuracy on CIFAR-10 and CIFAR-100 using the ResNet-18 (Base + Split (SP-2) + AWGN) model and the ResNet-34 (Base + Split (SP-2) + AWGN) model across different $z$ dimensions and SNR levels. Note that the names of the models above and the results on CIFAR-100 for SNR = 5 dB and $N_{c}=1024$ are consistent with those in Table I.
@@
TABLE I: Comparison of Top-1 Accuracy on CIFAR-100 under different settings, given SNR = 5 dB and $N_{c}$ = 1024. Gray shading denotes the same split point as shown in Fig. 2. indicates the identical models or models identical except for the split point.
@@
TABLE II: Comparison of FLOPs, Parameters, and Top-1 Accuracy on CIFAR-100 using ResNet-34 at different split points. Gray shading denotes the same split point as shown in Fig. 2. Both FLOPs and Parameters are measured in M (million).
@@
Since device capabilities are a critical consideration in practical networks, we calculate the FLOPs and parameter counts for each module. In the experiments reported in Fig. 3 and Table I, which use a single fractionally-strided convolution module, we observe that the decompression module at SP-2 is extremely costly, since its parameters and FLOPs increase by approximately 12.8 and 60 times, respectively. To address this, we adopt two fractionally-strided convolution (ConvTranspose2d) modules (as shown in Fig. 2), to achieve a more practical and balanced FLOP distribution between edge and cloud devices. The updated experimental results are presented in Fig. 4, covering four representative split locations after each residual layer block of ResNet-34 (SP-2 to SP-5). As expected, accuracy increases progressively from SP-2 to SP-5. Detailed numerical results on FLOPs, parameter distribution, and accuracy are summarized in Table II. At SP-2, the transmitter accounts for only about 10% of total computation. By SP-4, the transmitter’s share rises to 79.12% of FLOPs and 35.53% of parameters. At SP-5, nearly all computation (98.93%) is performed on the transmitter, leaving the receiver’s workload negligible. This evaluation helps identify an optimal split point that satisfies real-world hardware constraints while balancing edge-side efficiency and task accuracy. A more detailed analysis follows in the next subsection.
@@ IV-D Computation and Communication Cost
In practice, $\beta$ depends on the hardware. Edge devices (e.g., microcontrollers or mobile CPUs) are generally much slower than cloud servers or GPUs. FLOP latency can range from microseconds on edge hardware to picoseconds on cloud platforms, corresponding to a 4–5 order-of-magnitude difference.
We evaluate $T_{\text{comp}}$ for $\beta$ values ranging from $10^{-5}$ to 1. Figure 5 plots the normalized computation cost against $\log_{10}(\beta)$ for the selected split points.
@@
We also assess communication efficiency by analyzing the dimensionality of the semantic feature vector $z$ (i.e., $N_{c}$). Table III presents the minimum $N_{c}$ required to achieve a Top-1 Accuracy of 0.66, which corresponds to roughly 85% of the baseline ResNet-34 accuracy (Vanilla + Transforms, no splitting).
TABLE III: Minimum semantic feature size $N_{c}$ required to achieve 85% of baseline accuracy across different SNR levels.
SNR (dB) SP-2 SP-3 SP-4
0 1024 64 32
3 512 32 32
5 512 32 16
The results indicate that the Early split (SP-2) requires larger feature dimensions than Mid and Late splits to reach the same accuracy level. This reveals a core trade-off: early partitioning reduces computation at the transmitter but increases communication load.
Furthermore, as channel conditions degrade, a higher $N_{c}$ is needed to preserve task accuracy. This highlights the importance of dynamically adapting $N_{c}$ in response to changing channel quality, particularly in real-world deployments with changing wireless conditions.

### v2

@@ Abstract
Empowered by deep learning, semantic communication marks a paradigm shift from transmitting raw data to conveying task-relevant meaning, enabling more efficient and intelligent wireless systems. In this study, we explore a deep learning-based task-oriented communication framework that jointly considers classification performance, computational latency, and communication cost. We evaluate ResNets-based models on the CIFAR-10 and CIFAR-100 datasets to simulate real-world classification tasks in wireless environments. We partition the model at various points to simulate split inference across a wireless channel. By varying the split location and the size of the transmitted semantic feature vector, we systematically analyze the trade-offs between task accuracy and resource efficiency. Experimental results show that, with appropriate model partitioning and semantic feature compression, the system can retain over 85% of baseline accuracy while significantly reducing both computational load and communication overhead.
@@ I Introduction
Traditional wireless communication systems have been designed according to Shannon’s information theory, where the primary objective is to reliably transmit bits over noisy channels [12]. However, with the rapid rise of data-driven applications such as autonomous driving, real-time surveillance, and industrial automation, communication networks are now expected to support not only reliable data transfer but also intelligent, task-specific decision-making. This paradigm shift has prompted a rethinking of communication system design, leading to growing interest in semantic and task-oriented communication [1], which focuses on conveying information relevant to the underlying task, aligning more closely with how humans communicate efficiently and robustly in uncertain environments [9, 13, 2].
@@
Recent advances in machine learning, particularly deep learning, have significantly accelerated the development of the semantic communication systems. Deep neural networks (DNNs), such as convolutional neural networks (CNNs) for images or Transformers for language, can capture high-level abstractions. Semantic encoders and decoders can be co-trained with downstream inference tasks to extract and transmit only the most relevant features. For example, instead of sending the full-resolution image, the encoder may learn to transmit only task-critical semantic features (e.g., class labels, bounding boxes, or embeddings), greatly reducing communication cost without compromising task performance. This concept extends naturally to task-oriented communication [3], where performance is measured not by bit error rate but by how well a specific task (e.g., classification, detection, or control) is completed. Such systems are especially valuable in resource-constrained environments like wireless sensor networks or mobile edge computing, where devices have limited energy, bandwidth, and processing power.
@@
Despite recent progress, designing and deploying deep learning-based task-oriented communication systems remains challenging. These systems must integrate sensing, encoding, transmission, and inference into a unified pipeline, where model architectures and training objectives must align with system-level constraints such as latency, energy consumption, and bandwidth. While deep models can be trained offline using powerful infrastructure, real-time inference during deployment must be both accurate and efficient, requiring a careful balance between task performance and the use of computational and communication resources. Existing studies address only parts of this challenge [5, 10, 8, 11]. For instance, [5] introduced adaptive device–edge co-inference to better utilize distributed computing resources but did not consider communication cost and channel conditions. Later work in [10] focused on compressing intermediate features to reduce transmission load, yet treated model partitioning as a secondary concern. More recent research has examined the timeliness of task-oriented inference via Age of Information [8], but without investigating the optimal split decisions. Overall, a unified analysis of supervised learning models under different partitioning settings considering both channel noise and compression network complexity remains lacking.
@@
In this paper, we present a deep learning-based task-oriented communication framework that explicitly considers both task accuracy and latency. The latter includes computation delay (from model inference) and communication delay (from transmitting semantic features). We focus on classification tasks, adopting the 18-layer and 34-layer Residual Networks [4] as backbone models and evaluating performance on the CIFAR-10 and CIFAR-100 datasets [6]. Through extensive experiments, we investigate how different model partitioning strategies and semantic feature dimensions influence the trade-offs among accuracy, computational load, and communication cost, providing practical insights for the design of efficient task-oriented communication systems.
@@
The remainder of this paper is organized as follows. Sec. II introduces the system architecture. Sec. III presents the design and implementation of the classification system. Sec. IV evaluates the system under various configurations and analyzes the trade-offs between accuracy and latency. Sec. V concludes the paper and discusses future directions.
@@ II Task Latency in Deep Learning-Based Task-Oriented Communication
  • Transmitter-side inference: The classifier is fully deployed at the transmitter (edge device), and only the final classification result (label) is sent. This minimizes communication overhead but places high computational demands on resource-constrained edge devices.
  • Receiver-side inference: The full input is transmitted to the receiver (cloud device), which performs the classification. This reduces transmitter-side computation but incurs high communication cost.
@@
We illustrate the system architecture in Fig. 1. Let $M_{t}$ and $M_{r}$ denote the parts of the DNN model deployed at the transmitter and receiver, respectively. Their inference times are denoted as $T_{M_{t}}$ and $T_{M_{r}}$. In general, the inference time of a model depends on both its complexity and the hardware it runs on. We quantify model complexity using floating-point operations (FLOPs), denoted by $F_{M}$ for the model $M$, which provides a hardware-agnostic measure of computational demand. The actual computation time of the model $M$ can be estimated as a function of its FLOP counts, where the function is determined by the hardware settings of the device that runs the model. Inside the compute-bound regime, the function is linear and monotone, so we estimate the computation time as:
@@
Based on the equations (1) - (4), if the devices are fixed and communication settings, such as bandwidth and modulation scheme, are predetermined, then the selection of the model partitioning strategy and the size of the intermediate semantic representation $N_{c}$ jointly determine whether the task can be executed accurately and timely.
In the following sections, we design a task-oriented communication system for classification, and evaluate the trade-off between latency and accuracy under different model partitioning strategies and $N_{c}$ values.
@@ III System Design for Task-Oriented Classification
Figure 2: Illustration of ResNet-34 architecture and a split-inference task-oriented network based on the ResNet-34 Vanilla model (unmodified baseline version). Specifically, the first three parts (input, conv1, and conv2_x) are considered in the transmitter (encoder) as pre-processing, while the rest of them are considered in the receiver (decoder) as post-processing. The compression and decompression modules ensure the post-processing input is consistent with the pre-processing output while preserving invariance and equivariance to translation. The semantic representation with $z$ dimension is transmitted from the encoder to the decoder over the AWGN channel. The FLOPs and proportion of each module are listed below. (Zoomed in for better visualization.)
@@ III-A Deep Learning Model for Classification
We adopt the 18-layer and the 34-layer Residual Networks (ResNet-18 and ResNet-34) as the backbone models for our task-oriented communication system. These widely used CNN architectures are known for their robustness, modularity, and strong performance across a broad range of image classification tasks.
@@
Due to the widespread adoption and adaptability across datasets and domains, the ResNet-based architectures like ResNet-34 make our system design generalizable to a variety of classification tasks and application scenarios.
@@ III-C Wireless Channel Simulation with AWGN
where $\sigma^{2}$ denotes the noise variance and is controlled by the assigned signal-to-noise ratio (SNR) of the channel. The receiver obtains the noisy feature vector $\tilde{z}$ and restores it to its pre-compression size, then uses it as the input to $M_{r}$ to perform the remaining convolutional layers and finally the classification task. This setup enables us to assess the robustness and performance of the system under varying channel conditions. To ensure a seamless interface between the encoder and decoder, the output dimension of $M_{t}$ should match the input dimension of $M_{r}$, i.e., both must operate on a vector of size $N_{c}$. Training and evaluation are implemented in an end-to-end manner within the widely adopted deep learning framework, PyTorch [7]. The source code is available at the GitHub repository: https://github.com/sanyeungwang/Deep-Semantic-Inference-over-the-Air.
@@ IV-A Datasets
CIFAR-10: This dataset [6] contains 60,000 color images of size $32\times 32$ across 10 classes: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. Each class includes 6,000 images. The dataset is divided into five training batches and one test batch, each with 10,000 images. Each RGB image has a shape of $32\times 32\times 3$, and class labels are encoded as integers from 0 to 9.
CIFAR-100: Also introduced in [6], CIFAR-100 shares the same image format as CIFAR-10 but includes 100 classes with 600 images each. For every class, there are 500 training images and 100 test images, making the classification task more fine-grained. The dataset is split into a single training batch of 50,000 images and a test batch of 10,000 images.
@@ IV-C Comparisons With Different Settings
Based on the proposed task-oriented communication system, we first analyze performance under different $z$ dimensions and SNR levels, as shown in Fig. 3. We evaluate both ResNet-18 and ResNet-34, each split between conv_2 and conv_3 (referred to as SP-2), with one fractionally-strided convolution module (ConvTranspose2d) in the decompression module, and assuming unlimited computing power in the cloud device initially. Top-1 Accuracy refers to the comparison between the class with the highest probability and the ground-truth label. For CIFAR-10 shown in Figs. 3(a) and 3(b), we observe that higher SNR leads to better performance, but accuracy approaches 0.9 even for low-dimensional $z$. This saturation effect makes CIFAR-10 less suitable for analyzing the impact of model architecture and compression settings. In contrast, CIFAR-100 (Figs. 3(c) and 3(d)) achieves lower overall accuracy due to its finer-grained classes and smaller inter-class variance. The performance gap between different SNR levels is also more pronounced, highlighting its sensitivity to AWGN and its suitability for studying the effects of semantic compression and noise. Therefore, we focus on CIFAR-100 in subsequent experiments.
Figure 3: Comparison of Top-1 Accuracy on CIFAR-10 and CIFAR-100 using ResNet-18 (Base + Split (SP-2) + AWGN) and ResNet-34 (Base + Split (SP-2) + AWGN) across different $z$ dimensions (implemented by varying the number of channels $N_{c}$) and SNR levels. Note that the names of the models above and the results on CIFAR-100 for SNR = 5 dB and $N_{c}=1024$ are consistent with those in Table I.
@@
TABLE I: Comparison of Top-1 Accuracy on CIFAR-100 under different settings, given SNR = 5 dB and $N_{c}$ = 1024. Gray shading denotes the same split point in Fig. 2. indicates the identical models or models identical except for the split point.
@@
TABLE II: Comparison of FLOPs, Parameters, and Top-1 Accuracy on CIFAR-100 using ResNet-34 at different split points. Gray shading denotes the same split point in Fig. 2. Both FLOPs and Parameters are measured in M (million).
@@
Since device capabilities are a critical consideration in practical networks, we calculate the FLOPs and parameter counts for each module. In the experiments reported in Fig. 3 and Table I, which use a single fractionally-strided convolution module, we observe that the decompression module at SP-2 is extremely costly, since its parameters and FLOPs increase by approximately 12.8 and 60 times, respectively. To address this, we adopt two fractionally-strided convolution (ConvTranspose2d) modules (as shown in Fig. 2), to achieve a more practical and balanced FLOP distribution between edge and cloud devices. The updated experimental results are presented in Fig. 4, covering four representative split locations after each residual layer block of ResNet-34 (SP-2 to SP-5). As expected, accuracy increases progressively from SP-2 to SP-5. Detailed numerical results on FLOPs, parameter distribution, and accuracy are summarized in Table II. At SP-2, the transmitter accounts for only about 10% of total computation. By SP-4, the transmitter’s share rises to 79.12% of FLOPs and 35.53% of parameters. At SP-5, nearly all computation (98.93%) is performed on the transmitter, leaving the receiver’s workload negligible. This evaluation helps identify an optimal split point that satisfies real-world hardware constraints while balancing edge-side efficiency and task accuracy, which follows in the next subsection.
@@ IV-D Computation and Communication Cost
In practice, $\beta$ depends on the hardware. Edge devices (e.g., microcontrollers or mobile CPUs) are generally much slower than cloud servers or GPUs. FLOP latency can range from microseconds on edge hardware to picoseconds on cloud platforms, corresponding to a 4–5 order-of-magnitude difference. Therefore, we evaluate $T_{\text{comp}}$ for $\beta$ values ranging from $10^{-5}$ to 1. Figure 5 plots the normalized computation cost against $\log_{10}(\beta)$ for the selected split points.
@@
We evaluate communication efficiency by measuring the transmission time of the semantic feature vector $z$, normalized by the time required to send the original images. Figure 6 shows the normalized communication time for features of dimension $N_{c}$, where $N_{c}$ is the minimum size needed to achieve a Top-1 Accuracy of 0.66 (approximately 85% of the baseline ResNet-34 accuracy without model splitting).
Figure 6: Normalized communication time under different SNR levels.
The results show that all split configurations could significantly reduce communication cost compared with transmitting raw input data. However, the Early split (SP-2) requires higher overhead than Mid and Late splits to reach the same accuracy level. This reveals a core trade-off: early partitioning reduces computation at the transmitter but increases communication load. Furthermore, as channel conditions degrade, larger semantic features are needed to preserve accuracy. This highlights the importance of dynamically adapting the dimension of semantic feature, particularly in real-world deployments in changing wireless conditions.